An economy has a monetary base of 1,000 $1 bills. Calculate the money supply in scenarios. 

a. All money is held as demand deposits. Banks hold 100 percent of deposits as reserves. 

b. All money is held as demand deposits. Banks hold 20 percent of deposits as reserves. 

c. People hold equal amounts of currency and demand deposits. Banks hold 20 percent of deposits as reserves.
Answer #1

(a) money multiplier (m) = 1/reserve ratio.

m = 1 / 1 (note: reserve ratio = 100%)

Money multiplier (m) = 1

Money supply = money multiplier * monetary base

Money supply = (1) * ($1000).

Money supply = $1000

(b) money multiplier (m) = 1/ reserve ratio

m = 1/0.2 (note: reserve ratio is 20%).

Money multiplier = 5

Money supply = money multiplier * monetary base.

Money supply = (5) ($1000)

Money supply = $5000.

(c) people hold $500 as currency and $500 in deposit.

Currency deposit ratio (CD) = currency / deposit

CD = $500/$500.

CD = 1.

Reserve ratio (RR) = 0.2

Money multiplier (m) = (1+CD) / (CD + RR)

Money multiplier (m) = (1+1) / (1+0.2)

m = 2/1.2

m = 1.67

Money supply = money multiplier * monetary base.

Money supply = (1.67)*($1000)

Money supply = $1670